# Learning WENO for entropy stable schemes to solve conservation laws

### v1

@@ 4.2 Vertex Selection Algorithm
11:             $y_{*}^{(2)}=\frac{1}{8}(1+\psi^{-})-\gamma_{1}\psi^{-}$
12:             $x_{c}=\frac{1}{3}(2\gamma_{2}+x_{*})$,    $y_{c}=\frac{1}{3}(2\gamma_{1}+y_{*}^{(2)})$
13:             $\bm{\nu}=\left[(\gamma_{2},\gamma_{1}),(x_{*},\gamma_{1}),(\gamma_{2},y_{*}^{(2)}),(x_{c},y_{c}),(x_{c},y_{c})\right]$
14:         end if
15:    else
16:         $y_{*}=\frac{1}{8}(1+\psi^{-})-\gamma_{1}\psi^{-}$
17:         if $y_{*}<\gamma_{2}$ then
18:             $\bm{\nu}=\left[(\gamma_{2},\gamma_{1}),(\gamma_{1},\gamma_{1}),(\gamma_{2},\gamma_{2}),(\gamma_{1},\gamma_{2}),(0,0)\right]$
19:         else
20:             $x_{*}^{(2)}=\frac{1}{8}(1+\psi^{+})-\gamma_{2}\psi^{+}$
21:             $\bm{\nu}=\left[(\gamma_{2},\gamma_{1}),(\gamma_{1},\gamma_{1}),(\gamma_{2},\gamma_{2}),(\gamma_{1},y_{*}),(x_{*}^{(2)},\gamma_{2})\right]$
22:         end if
23:    end if
24:else if $\theta_{i+1}^{-}<1$, $\theta_{i}^{+}>1$ then
25:    if $\psi^{+}<-1$ then
26:         $x_{*}=\frac{1}{8}(1+\psi^{+})-\gamma_{1}\psi^{+}$
27:         if $x_{*}<\gamma_{2}$ then
28:             $\bm{\nu}=\left[(\gamma_{2},\gamma_{1}),(\gamma_{1},\gamma_{1}),(\gamma_{2},\gamma_{2}),(\gamma_{1},\gamma_{2}),(0,0)\right]$
29:         else
30:             $y_{*}^{(3)}=\frac{1}{8}(1+\psi^{-})-\gamma_{2}\psi^{-}$
31:             $\bm{\nu}=\left[(\gamma_{1},\gamma_{2}),(\gamma_{1},\gamma_{1}),(\gamma_{2},\gamma_{2}),(x_{*},\gamma_{1}),(\gamma_{2},y_{*}^{(3)})\right]$
32:         end if
33:    else
34:         $y_{*}=\frac{1}{8}(1+\psi^{-})-\gamma_{1}\psi^{-}$
35:         if $y_{*}<\gamma_{2}$ then
36:             $\hat{x}=\frac{(\psi^{-})^{2}+\psi^{-}}{8((\psi^{-})^{2}+1)}$,    $\hat{y}=\frac{1}{8}(1+\psi^{-})-\hat{x}\psi^{-}$
37:             $\bm{\nu}=\left[(\hat{x},\hat{y}),(\hat{x},\hat{y}),(\hat{x},\hat{y}),(\hat{x},\hat{y}),(\hat{x},\hat{y})\right]$
38:         else
39:             $x_{*}^{(3)}=\frac{1}{8}(1+\psi^{+})-\gamma_{1}\psi^{+}$
40:             $x_{c}=\frac{1}{3}(2\gamma_{1}+x_{*}^{(3)})$,    $y_{c}=\frac{1}{3}(2\gamma_{2}+y_{*})$
41:             $\bm{\nu}=\left[(\gamma_{1},\gamma_{2}),(x_{*}^{(3)},\gamma_{2}),(\gamma_{1},y_{*}),(x_{c},y_{c}),(x_{c},y_{c})\right]$
42:         end if
43:    end if
44:else if $\theta_{i+1}^{-}==1$, $\theta_{i}^{+}>1$ then
45:    $\bm{\nu}=\left[(\frac{1}{8},-\frac{3}{8}),(\frac{1}{8},\frac{1}{8}),(\frac{1}{8},\frac{1}{8}),(\frac{1}{8},-\frac{3}{8}),(\frac{1}{8},-\frac{1}{8})\right]$
46:else if $\theta_{i}^{+}==1$, $\theta_{i+1}^{-}>1$ then
47:    $\bm{\nu}=\left[(\frac{1}{8},\frac{1}{8}),(\frac{1}{8},\frac{1}{8}),(-\frac{3}{8},\frac{1}{8}),(-\frac{3}{8},\frac{1}{8}),(-\frac{1}{8},\frac{1}{8})\right]$
48:else
49:    $\bm{\nu}=\left[(\frac{1}{8},\frac{1}{8}),(\frac{1}{8},-\frac{3}{8}),(-\frac{3}{8},-\frac{3}{8}),(-\frac{3}{8},\frac{1}{8}),(-\frac{1}{8},-\frac{1}{8})\right]$
50:end if
4.3 Neural Network Architecture
Since the neural network only has to learn to generate five convex weights given a relatively small feature space as input, we use a fairly simple network architecture. In particular, we employ an MLP that consists of three hidden layers with five neurons each, and an output layer also containing five neurons, leading to an MLP with #$\#{\bm{\Phi}}=120$ trainable parameters. The ReLU activation function is used in the hidden layers. Following the output layer, we use the softmax activation function to ensure that the network predicts a vector of convex vertex weights $\bm{\alpha}$. The predicted $\bm{\alpha}$ are used to obtain the perturbation according to (34), which are then used to estimate the cell interface reconstructions according to (35).
Other variations of the the MLP’s architecture (including activation functions) were considered, but ultimately this architecture yielded the best results.
4.4 Training Setup
We use the $l_{1}$ loss function (36) to calculate the difference between the actual values at the cell interface and the reconstructed values that are obtained from DSP-WENO. The objective of the training process is to minimize this loss. To train the neural network, we use the Adam optimizer [25] with optimizer parameters $\beta_{1}=0.5$, $\beta_{2}=0.9$, and a weight decay regularization of $10^{-5}$. We use a learning rate of 0.001. We perform a training/validation/test split in a 0.6/0.2/0.2 proportion, so that we can assess the performance on withheld data (the validation set) and then test on unseen data (the test set) when the network is finished being trained. We use mini-batches of size 500 and reshuffle the training set every epoch. The network is trained using PyTorch for 50 epochs on a 2.60 GHz Intel Core i7-10750H CPU. The training time for training one instance of this MLP is less than 30 seconds. We perform five training runs with different random initializations of the network weight and biases and select the network that performs the best on the test set.
4.5 Properties of DSP-WENO
DSP-WENO by construction satisfies all properties of SP-WENO with the exception of the mirror property. This property is lost due to the nature of the vertex selection algorithm and computation of the convex weights by the neural network. While it is possible to have pursued a construction that preserves the mirror property, we choose to omit it so as not to overly constrain the network. Additionally, DSP-WENO satisfies the following stability estimate for the reconstructed jump in terms of the original jumps.
Lemma 3.
@@ 4.5 Properties of DSP-WENO
$\left|\llbracket z\rrbracket_{i+\frac{1}{2}}\right|\leq\frac{1}{2}\left|\Delta z_{i-\frac{1}{2}}\right|+\left|\Delta z_{i+\frac{1}{2}}\right|+\frac{1}{2}\left|\Delta z_{i+\frac{3}{2}}\right|\hskip 8.53581pt\forall\hskip 2.84526pti\in\mathbb{Z}.$ (39)
@@
If $\Delta z_{i+\frac{1}{2}}=0$, the bound will clearly be satisfied since $|\llbracket z\rrbracket_{i+\frac{1}{2}}|=0$ in this case by construction of DSP-WENO. Hence, we assume $\Delta z_{i+\frac{1}{2}}\neq 0$. We start with the explicit expression for $\llbracket z\rrbracket_{i+\frac{1}{2}}$ (17) and expand
$\llbracket z\rrbracket_{i+\frac{1}{2}}$ $=\frac{1}{2}\left[\tilde{w}_{0}\left(1-\theta_{i+1}^{-}\right)+w_{1}\left(1-\theta_{i}^{+}\right)\right]\Delta z_{i+\frac{1}{2}}$
$=\frac{1}{2}\left[\left(\frac{1}{4}-2C_{2}\right)\left(1-\frac{\Delta z_{i+\frac{3}{2}}}{\Delta z_{i+\frac{1}{2}}}\right)+\left(\frac{1}{4}-2C_{1}\right)\left(1-\frac{\Delta z_{i-\frac{1}{2}}}{\Delta z_{i+\frac{1}{2}}}\right)\right]\Delta z_{i+\frac{1}{2}}$
$=\left(C_{1}-\frac{1}{8}\right)\Delta z_{i-\frac{1}{2}}+\left(\frac{1}{4}-C_{1}-C_{2}\right)\Delta z_{i+\frac{1}{2}}+\left(C_{2}-\frac{1}{8}\right)\Delta z_{i+\frac{3}{2}}.$
By consistency, we have $-3/8\leq C_{1},C_{2}\leq 1/8$. Thus, the absolute value of the jump is clearly bounded
@@
Remark 9.
The estimate (39) is not unique to DSP-WENO and is a consequence of the sign property constraint (18). Thus, it is also satisfied by every variant of SP-WENO. However, it was shown in [16] that the original SP-WENO satisfies a sharper bound which is attributed to the fact the the reconstructed jump is zero is many cases.
5 Numerical Results
In this section, we present several numerical results to demonstrate the efficacy of the proposed DSP-WENO approach. We consider both 1D and 2D test cases for the evolution of scalar and systems of conservation laws. In 1D, we discretize a domain $[a,b]$ using an $N$-cell uniform mesh with $h=\frac{b-a}{N}$
$x_{i+\frac{1}{2}}:=a+ih,\hskip 8.53581ptx_{i}=\frac{x_{i-\frac{1}{2}}+x_{i+\frac{1}{2}}}{2},\hskip 8.53581pti=0,...,N.$
In 2D, we discretize the domain $[a,b]\times[c,d]$ using a uniform mesh (in each dimension) with $h_{x}=\frac{b-a}{N_{x}}$ and $h_{y}=\frac{d-c}{N_{y}}$. The mesh is constructed as
$x_{i+\frac{1}{2}}:=a+ih_{x},\hskip 8.53581ptx_{i}=\frac{x_{i-\frac{1}{2}}+x_{i+\frac{1}{2}}}{2},\hskip 8.53581pti=0,...,N_{x},$
$y_{j+\frac{1}{2}}:=c+jh_{y},\hskip 8.53581pty_{j}=\frac{y_{j-\frac{1}{2}}+y_{j+\frac{1}{2}}}{2},\hskip 8.53581ptj=0,...,N_{y}.$
Ghost cells are introduced to extend the mesh as needed while imposing either periodic or Neumann boundary conditions. In all evolution cases, we use the TeCNO4 numerical flux, consisting of the fourth-order accurate entropy conservative flux (7) and various sign-preserving reconstruction methods for the diffusion operator. The CFL may vary across problems and is specified in each case. SSP-RK3 (see [18]) is used for time-marching in all evolution test cases.
5.1 Reconstruction Accuracy
Before proceeding to solve conservation laws, we demonstrate the reconstruction accuracy of DSP-WENO. We consider the reconstruction of a smooth inclined sine wave
$u(x)=\sin{(10\pi x)}+x,\hskip 8.53581ptx\in[0,1],$
as the mesh size $h$ is varied. To compute the error in the interface values, we evaluate
$\|u^{-}_{i+\frac{1}{2}}-u(x_{i+\frac{1}{2}})\|_{L^{1}_{h}}+\|u^{+}_{i+\frac{1}{2}}-u(x_{i+\frac{1}{2}})\|_{L^{1}_{h}},$
where the discrete $L^{1}$ norm is defined as
$\|(\cdot)_{i}\|_{L^{1}_{h}}=\sum_{i=1}^{N}|(\cdot)_{i}|h.$
Table 3 shows the errors and corresponding convergence rates for ENO3, SP-WENO, SP-WENOc, and DSP-WENO. All reconstruction methods achieve the expected third-order convergence with the SP-WENO variants exhibiting super-convergence. Note that the errors of DSP-WENO are about an order of magnitude larger than those of SP-WENO and SP-WENOc while being similar to ENO3 on the finer meshes.
Now, we consider a continuous function with a sharp connection between two linear components to mimic a discontinuity:
@@ 5.1 Reconstruction Accuracy
where $\epsilon$ is chosen as 0.001. We reconstruct on a mesh consisting of 1000 cells, ensuring that a single cell is sitting inside the jump. In other words, we are simulating a scenario where the discontinuity is not perfectly resolved. Figure 4 shows the left and right reconstructed interface values for ENO3, SP-WENO, SP-WENOc, and DSP-WENO, zoomed into the region close to the simulated discontinuity. Figure 4 also features bar plots that show the actual jump in cell center values and the reconstructed jump at the interface for each method. The left and right reconstructed values are almost identical outside of the $0\leq x\leq\epsilon$ region for all the methods, and thus not visible in the figure given the scale of the jump. However, inside the pseudo-shock where a cell (let us call it $I_{0}$) lies, SP-WENO and SP-WENOc do not fully capture the non-zero jumps. At the left interface of $I_{0}$, SP-WENO results in a very small reconstructed jump, while SP-WENOc leads to a larger jump but which is still an order of magnitude smaller as compared to the jump at the right interface of $I_{0}$. On the other hand, ENO3 and DSP-WENO give rise to larger reconstructed jumps at both interfaces of $I_{0}$, which is the desirable behavior in the vicinity of discontinuities. In the numerical results that follow, we will see a visible impact of the reconstructed jump magnitude.
N ENO3 SP-WENO SP-WENOc DSP-WENO
$L_{h}^{1}$ $L_{h}^{1}$ $L_{h}^{1}$ $L_{h}^{1}$
Error Rate Error Rate Error Rate Error Rate
40 3.47e-2 - 7.27e-2 - 7.41e-2 - 1.65e-1 -
80 4.54e-3 2.93 5.85e-3 3.64 6.37e-3 3.54 3.01e-2 2.45
160 5.84e-4 2.96 4.45e-4 3.72 4.71e-4 3.76 2.83e-3 3.41
320 7.42e-5 2.98 3.29e-5 3.76 3.43e-5 3.78 2.14e-4 3.73
640 9.38e-6 2.98 2.37e-6 3.79 2.46e-6 3.80 1.55e-5 3.78
1280 1.17e-6 3.00 1.68e-7 3.82 1.74e-7 3.82 1.22e-6 3.67
2560 1.47e-7 3.00 1.18e-8 3.84 1.21e-8 3.84 1.13e-7 3.44
Table 3: Reconstruction accuracy for inclined sine wave.

### v2

@@ 1 $\theta_{i}^{+},\theta_{i+1}^{-}>1$ $C_{1}=C_{2}=\frac{1}{8}$ None $\tilde{w}_{0}=w_{1}=0$
11:             $y_{*}^{(2)}=\frac{1}{8}(1+\psi^{-})-\gamma_{1}\psi^{-}$,   $x_{c}=\frac{1}{3}(2\gamma_{2}+x_{*})$,    $y_{c}=\frac{1}{3}(2\gamma_{1}+y_{*}^{(2)})$
12:             $\bm{\nu}=\left[(\gamma_{2},\gamma_{1}),(x_{*},\gamma_{1}),(\gamma_{2},y_{*}^{(2)}),(x_{c},y_{c}),(x_{c},y_{c})\right]$
13:         end if
14:    else
15:         $y_{*}=\frac{1}{8}(1+\psi^{-})-\gamma_{1}\psi^{-}$,    $x_{*}^{(2)}=\frac{1}{8}(1+\psi^{+})-\gamma_{2}\psi^{+}$
16:         if $y_{*}<\gamma_{2}$ then
17:             $\bm{\nu}=\left[(\gamma_{2},\gamma_{1}),(\gamma_{1},\gamma_{1}),(\gamma_{2},\gamma_{2}),(\gamma_{1},\gamma_{2}),(0,0)\right]$
18:         else
19:             $\bm{\nu}=\left[(\gamma_{2},\gamma_{1}),(\gamma_{1},\gamma_{1}),(\gamma_{2},\gamma_{2}),(\gamma_{1},y_{*}),(x_{*}^{(2)},\gamma_{2})\right]$
20:         end if
21:    end if
22:else if $\theta_{i+1}^{-}<1$, $\theta_{i}^{+}>1$ then
23:    if $\psi^{+}<-1$ then
24:         $x_{*}=\frac{1}{8}(1+\psi^{+})-\gamma_{1}\psi^{+}$,    $y_{*}^{(3)}=\frac{1}{8}(1+\psi^{-})-\gamma_{2}\psi^{-}$
25:         if $x_{*}<\gamma_{2}$ then
26:             $\bm{\nu}=\left[(\gamma_{2},\gamma_{1}),(\gamma_{1},\gamma_{1}),(\gamma_{2},\gamma_{2}),(\gamma_{1},\gamma_{2}),(0,0)\right]$
27:         else
28:             $\bm{\nu}=\left[(\gamma_{1},\gamma_{2}),(\gamma_{1},\gamma_{1}),(\gamma_{2},\gamma_{2}),(x_{*},\gamma_{1}),(\gamma_{2},y_{*}^{(3)})\right]$
29:         end if
30:    else
31:         $y_{*}=\frac{1}{8}(1+\psi^{-})-\gamma_{1}\psi^{-}$
32:         if $y_{*}<\gamma_{2}$ then
33:             $\hat{x}=\frac{(\psi^{-})^{2}+\psi^{-}}{8((\psi^{-})^{2}+1)}$,    $\hat{y}=\frac{1}{8}(1+\psi^{-})-\hat{x}\psi^{-}$
34:             $\bm{\nu}=\left[(\hat{x},\hat{y}),(\hat{x},\hat{y}),(\hat{x},\hat{y}),(\hat{x},\hat{y}),(\hat{x},\hat{y})\right]$
35:         else
36:             $x_{*}^{(3)}=\frac{1}{8}(1+\psi^{+})-\gamma_{1}\psi^{+}$,    $x_{c}=\frac{1}{3}(2\gamma_{1}+x_{*}^{(3)})$,    $y_{c}=\frac{1}{3}(2\gamma_{2}+y_{*})$
37:             $\bm{\nu}=\left[(\gamma_{1},\gamma_{2}),(x_{*}^{(3)},\gamma_{2}),(\gamma_{1},y_{*}),(x_{c},y_{c}),(x_{c},y_{c})\right]$
38:         end if
39:    end if
40:else if $\theta_{i+1}^{-}==1$, $\theta_{i}^{+}>1$ then
41:    $\bm{\nu}=\left[(\frac{1}{8},-\frac{3}{8}),(\frac{1}{8},\frac{1}{8}),(\frac{1}{8},\frac{1}{8}),(\frac{1}{8},-\frac{3}{8}),(\frac{1}{8},-\frac{1}{8})\right]$
42:else if $\theta_{i}^{+}==1$, $\theta_{i+1}^{-}>1$ then
43:    $\bm{\nu}=\left[(\frac{1}{8},\frac{1}{8}),(\frac{1}{8},\frac{1}{8}),(-\frac{3}{8},\frac{1}{8}),(-\frac{3}{8},\frac{1}{8}),(-\frac{1}{8},\frac{1}{8})\right]$
44:else
45:    $\bm{\nu}=\left[(\frac{1}{8},\frac{1}{8}),(\frac{1}{8},-\frac{3}{8}),(-\frac{3}{8},-\frac{3}{8}),(-\frac{3}{8},\frac{1}{8}),(-\frac{1}{8},-\frac{1}{8})\right]$
46:end if
Case 1
Case 2
Case 3
Case 4a
Case 5a
Cases 4b, 5b, 6
SM3 Stability bound for DSP-WENO
Lemma SM1.
@@
$\left|\llbracket z\rrbracket_{i+\frac{1}{2}}\right|\leq\frac{1}{2}\left|\Delta z_{i-\frac{1}{2}}\right|+\left|\Delta z_{i+\frac{1}{2}}\right|+\frac{1}{2}\left|\Delta z_{i+\frac{3}{2}}\right|\hskip 8.53581pt\forall\hskip 2.84526pti\in\mathbb{Z}.$ (SM3)
@@
If $\Delta z_{i+\frac{1}{2}}=0$, the bound will clearly be satisfied since $|\llbracket z\rrbracket_{i+\frac{1}{2}}|=0$ in this case by construction of DSP-WENO. Hence, we assume $\Delta z_{i+\frac{1}{2}}\neq 0$. We start by using (13) to rewrite the expression of $\llbracket z\rrbracket_{i+\frac{1}{2}}$ in (15) as
$\llbracket z\rrbracket_{i+\frac{1}{2}}=\left(C_{1}-\frac{1}{8}\right)\Delta z_{i-\frac{1}{2}}+\left(\frac{1}{4}-C_{1}-C_{2}\right)\Delta z_{i+\frac{1}{2}}+\left(C_{2}-\frac{1}{8}\right)\Delta z_{i+\frac{3}{2}}.$
By consistency, we have $-3/8\leq C_{1},C_{2}\leq 1/8$. Thus, the absolute value of the jump can be bounded as
@@
Remark SM5.
The estimate (SM3) is not unique to DSP-WENO and is a consequence of the sign property constraint (16). Thus, it is also satisfied by every variant of SP-WENO. However, the original SP-WENO satisfies a sharper bound which is attributed to the fact that the reconstructed jump is zero in most cases [fjordholm2016sign].
SM4 Additional Numerical Results and Details
SM4.1 Empirical study of reconstruction jumps
We consider a continuous function with a sharp connection between two linear components to mimic a discontinuity:
@@
where $\epsilon$ is chosen as 0.001. We reconstruct on a mesh consisting of 1000 cells, ensuring that a single cell is sitting inside the jump. In other words, we are simulating a scenario where the discontinuity is not perfectly resolved. Figure SM2 shows the left and right reconstructed interface values for ENO3, SP-WENO, SP-WENOc, and DSP-WENO, zoomed into the region close to the simulated discontinuity. Figure SM2 also features bar plots that show the actual jump in cell center values and the reconstructed jump at the interface for each method. The left and right reconstructed values are almost identical outside of the $0\leq x\leq\epsilon$ region for all the methods, and thus not visible in the figure given the scale of the jump. However, inside the pseudo-shock where a cell (let us call it $I_{0}$) lies, SP-WENO and SP-WENOc do not fully capture the non-zero jumps. At the left interface of $I_{0}$, SP-WENO results in a very small reconstructed jump, while SP-WENOc leads to a larger jump but which is still an order of magnitude smaller as compared to the jump at the right interface of $I_{0}$. On the other hand, ENO3 and DSP-WENO give rise to larger reconstructed jumps at both interfaces of $I_{0}$, which is the desirable behavior in the vicinity of discontinuities. The impact of the reconstructed jump magnitude can be observed from the numerical results for conservation laws shown in Section 5 in the main text.